OpenStudy (anonymous):

-7/3 - 5/6 + (2/3 ^2)

OpenStudy (pfenn1):

Is this the expression?\[\frac{-7}{3} - \frac56 + (\frac23)^2\]

OpenStudy (anonymous):

yea

OpenStudy (pfenn1):

You need to find the LCD of the three terms. What would it be?

OpenStudy (anonymous):

whats the LCD?

OpenStudy (pfenn1):

Lowest Common Denominator

OpenStudy (anonymous):

i have no idea

OpenStudy (pfenn1):

We want to be able to add and subtract the fractions, but we can't because the denominators are all different. So we want to come up with one that we can translate all the fractions with. The denominators are 3,6=(2)(3), and 9=(3)(3). So let's try(2)(3)(3)=18. So let's use that\[=\frac{-7}{3} - \frac56 + \frac49=\frac{-42-15+8}{18}=?\]

OpenStudy (anonymous):

-49/18

OpenStudy (anonymous):

??

OpenStudy (pfenn1):

That doesn't sound right to you?